OpenStudy (anonymous):

Will give medal!! Find the slope of the line that passes through each pair of points Question 3: A (-2, 1), B (4, 2) Question 3: A (5, 3), B (-1, 2)

OpenStudy (bradely):

A (-2, 1), B (4, 2) m = (y2 -y1)/(x2 -x1) =(2-1)/(4-(-2)) =1/6 For any help post on http://www.mathskey.com/